OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. A(5, 5), m = 3

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@TheMysteriousRainbowPegasus

OpenStudy (kaylieanime):

Do you know the point slope form?

OpenStudy (anonymous):

nope

OpenStudy (kaylieanime):

Well it's \[y - y _{1} = m(x - x _{1})\]

OpenStudy (anonymous):

what do i put where?

OpenStudy (anonymous):

I'm not the best at math. Sorry. :(

OpenStudy (kaylieanime):

I'd say y - 5 = 3(x - 5). Can you simplify?

OpenStudy (kaylieanime):

Maybe it's better if I guide... the equation above simplifies to y - 5 = 3x - 15. Add 5 to both sides: y = 3x - 10.

OpenStudy (kaylieanime):

And there's your slope intercept form.

OpenStudy (anonymous):

y = 8x - 15?

OpenStudy (kaylieanime):

(No, I gave the answer slightly) but since we need the standard form, we need to move 3x to the other side, making it look like this: y - 3x = 10.

OpenStudy (anonymous):

ok thnx